# Multispecies inhomogeneous $t$-PushTASEP from antisymmetric fusion

### v2

@@ 3.1. Space $V^{k}$ with base labeled with $\mathscr{B}^{k}$ and $\mathscr{T}^{k}$
For $10\leq k\leq n+$ 1, set
$\mathscr{B}^{k}$ =$=\{{\bf i}=(i_{0},\ldots,i_{n})\in\{0,1\}^{n+1}\mid|{\bf i}|=k\},\quad(|{\bf i}|=i_{0}+\cdots+i_{n}),$ (13)
$V^{k}$ $=\bigoplus_{{\bf i}\in\mathscr{B}^{k}}\mathbb{C}v_{\bf i}.$ (14)
One has $\dim V^{k}=\binom{n+1}{k}$. For the special case $k=1$, we identify $V^{1}$ with ${\mathsf{V}}$, the space of local states of the $t$-PushTASEP, via [^4] In Appendix A, $v_{\bf i}\in V^{k}$ with general $k$ will be identified with the antisymmetric tensor rather than the simple monomial ${\mathsf{v}}_{i_{1}}\otimes\cdots\otimes{\mathsf{v}}_{i_{k}}\in{\mathsf{V}}^{\otimes k}$. However such a connection is used only to explain the fusion procedure there.
${\mathsf{V}}\ni{\mathsf{v}}_{i}=v_{{\bf e}_{i}}\in V^{1}\;\;\text{where}\;\;{\bf e}_{i}=(\delta_{0,i},\ldots,\delta_{n,i})\in\mathscr{B}^{1}\quad(0\leq i\leq n).$ (15)
@@
$\mathscr{T}^{k}=\{{\bf I}=(I_{1},\ldots,I_{k})\mid 0\leq I_{1}<\cdots<I_{k}\leq n\},$ (16)
@@
We identify $\mathscr{B}^{k}$ and $\mathscr{T}^{k}$ by the one-to-one correspondence where $i_{\alpha}=0,1$ in ${\bf i}\in\mathscr{B}^{k}$ is regarded as the multiplicity of the letter $\alpha$ in ${\bf I}\in\mathscr{T}^{k}$. The arrays ${\bf i}$ in (13) and ${\bf I}\in\mathscr{T}^{k}$ in (16) will be referred to as the multiplicity representation and the tableau representation, respectively.
3.2. 3D construction of $R(z)$
$\mathcal{L}^{0,0}_{0,0}=\mathcal{L}^{1,1}_{1,1}=1,\quad\mathcal{L}^{1,0}_{1,0}={\rm{\bf k}},\quad\mathcal{L}^{0,1}_{0,1}=-q{\rm{\bf k}},\quad\mathcal{L}^{1,0}_{0,1}={\rm{\bf a}}^{+},\quad\mathcal{L}^{0,1}_{1,0}={\rm{\bf a}}^{-},$ (17a)
$\mathcal{L}^{a,b}_{i,j}=0\quad\text{if}\;\;a+b\neq i+j.$ (17b)
Here ${\rm{\bf k}},{\rm\mathbf{a}}^{\!+},{\rm\mathbf{a}}^{\!-}$ are $q$-oscillator operators [^5] The parameter $q$ will be set to $t^{1/2}$ in (26). on the Fock space $F=\bigoplus_{m\geq 0}\mathbb{Q}(q)|m\rangle$, defined by
${\rm{\bf k}}|m\rangle=q^{m}|m\rangle,\quad{\rm\mathbf{a}}^{\!+}|m\rangle=|m+1\rangle,\quad{\rm\mathbf{a}}^{\!-}|m\rangle=(1-q^{2m})|m-1\rangle.$ (18)
@@
Figure 1. $\mathcal{L}=(\mathcal{L}^{a,b}_{i,j})$ as a $q$-oscillator-weighted six-vertex model. The $q$-oscillators may be regarded as acting along the third arrow perpendicular to the sheet in each vertex. In this context, $\mathcal{L}$ is referred to as a 3D $L$-operator.
For $10\leq k,l\leq n+$ 1, we introduce the linear map $R(z)=R^{k,l}(z)\in\mathrm{End}(V^{k}\otimes V^{l})$ by
$R(z)(v_{\bf i}\otimes v_{\bf j})$ $=\sum_{{\bf a}\in\mathscr{B}^{k},{\bf b}\in\mathscr{B}^{l}}R(z)^{{\bf a},{\bf b}}_{{\bf i},{\bf j}}\,v_{\bf a}\otimes v_{\bf b}\qquad({\bf i}\in\mathscr{B}^{k},{\bf j}\in\mathscr{B}^{l}),$ (19a)
$R(z)^{{\bf a},{\bf b}}_{{\bf i},{\bf j}}$ $=\mathrm{Tr}_{F}(z^{\bf h}\mathcal{L}^{a_{n},b_{n}}_{i_{n},j_{n}}\cdots\mathcal{L}^{a_{0},b_{0}}_{i_{0},j_{0}}).$ (19b)
The trace is convergent as a formal power series in $q$ and $z$. From (17b), $R(z)$ has the weight conservation property:
$R(z)^{{\bf a},{\bf b}}_{{\bf i},{\bf j}}=0\;\,\text{unless}\;\,{\bf a}+{\bf b}={\bf i}+{\bf j}.$ (20)
The cases $k=0,n+1$ reduce to the scalar matrices as
$R^{0,l}(z)$ $=\mathrm{Tr}_{F}\left(z^{\bf h}(\mathcal{L}^{0,0}_{0,0})^{n+1-l}(\mathcal{L}^{0,1}_{0,1})^{l}\right)\mathrm{Id}=\frac{(-q)^{l}}{1-q^{l}z}\mathrm{Id},$ (21)
$R^{n+1,l}(z)$ $=\mathrm{Tr}_{F}\left(z^{\bf h}(\mathcal{L}^{1,1}_{1,1})^{l}(\mathcal{L}^{1,0}_{1,0})^{n+1-l}\right)\mathrm{Id}=\frac{1}{1-q^{n+1-l}z}\mathrm{Id}.$ (22)
The first nontrivial case is $l=1$. We express the elements ${\bf b},{\bf j}\in\mathscr{B}^{1}$ as ${\bf e}_{b},{\bf e}_{j}$ with $0\leq b,j\leq n$. Then (19b) is evaluated explicitly as
$(1-q^{k-1}z)(1-q^{k+1}z)R(z)^{{\bf a},{\bf e}_{b}}_{{\bf i},{\bf e}_{j}}$ $=\delta^{{\bf a}+{\bf e}_{b}}_{{\bf i}+{\bf e}_{j}}\times\begin{cases}(-q)^{1-a_{j}}(1-q^{2a_{j}+k-1}z),&j=b,\\ q^{a_{j+1}\cdots+a_{b-1}}(1-q^{2}),&j<b,\\ q^{k-1-(a_{b+1}+\cdots+a_{j-1})}(1-q^{2})z,&j>b.\end{cases}$ (23)
$R^{k_{1},k_{2}}_{1,2}(x)R^{k_{1},k_{3}}_{1,3}(xy)R^{k_{2},k_{3}}_{2,3}(y)=R^{k_{2},k_{3}}_{2,3}(y)R^{k_{1},k_{3}}_{1,3}(xy)R^{k_{1},k_{2}}_{1,2}(x)\qquad(k_{1},k_{2},k_{3}\in\{0,\ldots,n+1\}).$ (24)
They are equalities in $\mathrm{End}(V^{k_{1}}\otimes V^{k_{2}}\otimes V^{k_{3}})$ on which $R^{k_{i},k_{j}}_{i,j}(z)$ acts on the $i$’th and the $j$’th components as $R^{k_{i},k_{j}}(z)$ and identity elsewhere. Details can be found in [K22, Chap. 11].
3.3. Modifying $R(z)$ into $S(z)$
${\tilde{S}}(z)(v_{\bf i}\otimes v_{{\bf j}})$ $=\sum_{{\bf a}\in\mathscr{B}^{k},{\bf b}\in\mathscr{B}^{l}}{\tilde{S}}(z)^{{\bf a},{\bf b}}_{{\bf i},{\bf j}}\,v_{\bf a}\otimes v_{{\bf b}}\qquad({\bf i}\in\mathscr{B}^{k},{\bf j}\in\mathscr{B}^{l}),$ (25)
${\tilde{S}}(z)^{{\bf a},{\bf b}}_{{\bf i},{\bf j}}$ $=(-q)^{k-l+\eta}R(q^{l-k}z)^{{\bf a},{\bf b}}_{{\bf i},{\bf j}}\left|{}_{q\rightarrow t^{1/2}}\right.,$ (26)
$\eta$ $=\eta^{{\bf a},{\bf b}}_{{\bf i},{\bf j}}=\sum_{0\leq r<s\leq n}(b_{r}a_{s}-i_{r}j_{s}).$ (27)
The quantity $\eta$ (27) is formally the same as [KMMO16, eq. (16)]. Obviously, ${\tilde{S}}(z)$ also possesses the weight conservation property as (20). Moreover, the Yang-Baxter equation (24) for $R^{k,l}(z)$ and the same argument as in the proof of [KMMO16, Prop.4] imply that ${\tilde{S}}^{k,l}(z)$ also satisfies the Yang-Baxter equation:
${\tilde{S}}^{k_{1},k_{2}}_{1,2}(x){\tilde{S}}^{k_{1},k_{3}}_{1,3}(xy){\tilde{S}}^{k_{2},k_{3}}_{2,3}(y)={\tilde{S}}^{k_{2},k_{3}}_{2,3}(y){\tilde{S}}^{k_{1},k_{3}}_{1,3}(xy){\tilde{S}}^{k_{1},k_{2}}_{1,2}(x)\qquad(k_{1},k_{2},k_{3}\in\{0,\ldots,n+1\}).$ (28)
$S(z)(v_{\bf i}\otimes v_{{\bf e}_{j}})$ $=\sum_{{\bf a}\in\mathscr{B}^{k},{\bf e}_{b}\in\mathscr{B}^{1}}S(z)^{{\bf a},{\bf e}_{b}}_{{\bf i},{\bf e}_{j}}\,v_{\bf a}\otimes v_{{\bf e}_{b}}\qquad({\bf i}\in\mathscr{B}^{k},{\bf e}_{j}\in\mathscr{B}^{1}),$ (29)
$S(z)^{{\bf a},{\bf e}_{b}}_{{\bf i},{\bf e}_{j}}$ $=(1-z)(1-q^{2}z)(-q)^{k-1+\eta}R(q^{1-k}z)^{{\bf a},{\bf e}_{b}}_{{\bf i},{\bf e}_{j}}\left|{}_{q\rightarrow t^{1/2}}\right.,$ (30)
$\eta$ $=\eta^{{\bf a},{\bf e}_{b}}_{{\bf i},{\bf e}_{j}}=\sum_{s>b}a_{s}-\sum_{r<j}i_{r},$ (31)
where, from (20), we assume ${\bf a}+{\bf e}_{b}={\bf i}+{\bf e}_{j}$. This leads to the expression
$k-1+\eta=\begin{cases}a_{j}-1+2(a_{j+1}+\cdots+a_{n}),&j=b,\\ -(a_{j+1}+\cdots+a_{b-1})+2(a_{j+1}+\cdots+a_{n}),&j<b,\\ (a_{b+1}+\cdots+a_{j-1})+2(a_{j+1}+\cdots+a_{n}),&j>b.\end{cases}$ (32)
Therefore the matrix element (30) takes the form
$S(z)^{{\bf a},{\bf e}_{b}}_{{\bf i},{\bf e}_{j}}$ $=\delta^{{\bf a}+{\bf e}_{b}}_{{\bf i}+{\bf e}_{j}}(-1)^{a_{0}+\cdots+a_{j-1}+i_{0}+\cdots+i_{b-1}}t^{a_{j+1}+\cdots+a_{n}}(1-t^{a_{j}}z^{\delta_{b,j}})z^{[j>b]}.$ (33)
The sign factor in (33) can also be expressed as $(-1)^{a_{r+1}+\cdots+a_{s-1}}$ with $r=\min(j,b)$ and $s=\max(j,b)$. It is noteworthy that (33) is a polynomial in both $t$ and $z$.
$S(z)^{{\bf e}_{i},{\bf e}_{i}}_{{\bf e}_{i},{\bf e}_{i}}=1-tz,\quad S(z)^{{\bf e}_{i},{\bf e}_{j}}_{{\bf e}_{i},{\bf e}_{j}}=(1-z)t^{[i>j]}\;(i\neq j),\quad S(z)^{{\bf e}_{j},{\bf e}_{i}}_{{\bf e}_{i},{\bf e}_{j}}=(1-t)z^{[i<j]}\;(i\neq j).$ (34)
The elements (34) are positive in some range of $t,z$, and the sum $\sum_{0\leq a,b\leq n}S(z)^{{\bf e}_{a},{\bf e}_{b}}_{{\bf e}_{i},{\bf e}_{j}}=1-tz$ is independent of $i,j$. It is well known that these properties can be utilized to construct a Markov Matrix of multispecies ASEP. See Section 7.
On the other hand for $k\neq 1$ in general, (33) is neither positive definite nor negative definite for fixed $t$ and $z$. Furthermore, the summation $\sum_{{\bf a}\in\mathscr{B}^{k},{\bf e}_{b}\in\mathscr{B}^{1}}S(z)^{{\bf a},{\bf e}_{b}}_{{\bf i},{\bf e}_{j}}$ does not become independent of ${\bf i},{\bf e}_{j}$. Consequently, the $R$-matrix $S(z)$ is not stochastic in the sense of [KMMO16]. What is intriguing, as revealed by our subsequent analysis, is that the Markov Matrix of the multispecies $t$-PushTASEP is nonetheless reproduced as a suitable linear combination of the transfer matrices constructed from $S^{0,1}(z),S^{1,1}(z),\ldots,S^{n+1,1}(z)$.
We depict the element (33) of $S^{k,1}(z)$ as
${S(z)^{{\bf a},{\bf e}_{b}}_{{\bf i},{\bf e}_{j}}}=$$j$$b$${\bf i}$${\bf a}$$z$
@@ 3.3. Modifying $R(z)$ into $S(z)$
Thus far, we have derived the $R$-matrices $S^{0,1}(z),S^{1,1}(z),\ldots,S^{n+1,1}(z)$ based on $\mathcal{L}$ by invoking the so-called 3D construction. An alternative approach to constructing them is the fusion procedure [KRS81] starting from the basic one $S^{1,1}(z)$ given in (33). A slight peculiarity in this case is that the fusion must be carried out using the degeneracy of the $R$-matrix corresponding to the antisymmetric tensor, as opposed to the symmetric tensor commonly used in much of the existing literature on integrable probability. Further details are provided in Appendix A. We note that the matrix elements (23) have essentially appeared as the basic ingredient in the vertex operator approach in [DO94].

### v3

@@ 3.1. Space $V^{k}$ with base labeled with $\mathscr{B}^{k}$ and $\mathscr{T}^{k}$
For $0\leq k\leq n+1$, set
$\mathscr{B}^{k}$ $=\{{\bf i}=(i_{0},\ldots,i_{n})\in\{0,1\}^{n+1}\mid|{\bf i}|=k\},\quad(|{\bf i}|=i_{0}+\cdots+i_{n}),$ (3.1)
$V^{k}$ $=\bigoplus_{{\bf i}\in\mathscr{B}^{k}}\mathbb{C}v_{\bf i}.$ (3.2)
One has $\dim V^{k}=\binom{n+1}{k}$. For the special case $k=1$, we identify $V^{1}$ with $\mathsf{V}$}, the space of local states of the $t$-PushTASEP, via [^4] In Section 3.2, $v_{\bf i}\in V^{k}$ with general $k$ will be identified with the antisymmetric tensor rather than the simple monomial ${\mathsf{v}}_{i_{1}}\otimes\cdots\otimes{\mathsf{v}}_{i_{k}}\in{\mathsf{V}}^{\otimes k}$. However such a connection is used only to explain the fusion procedure there.
${\mathsf{V}}\ni{\mathsf{v}}_{i}=v_{{\bf e}_{i}}\in V^{1}\;\;\text{where}\;\;{\bf e}_{i}=(\delta_{0,i},\ldots,\delta_{n,i})\in\mathscr{B}^{1}\quad(0\leq i\leq n).$ (3.3)
@@
$\mathscr{T}^{k}=\{{\bf I}=(I_{1},\ldots,I_{k})\mid 0\leq I_{1}<\cdots<I_{k}\leq n\},$ (3.4)
@@
0
1
0
2
0
3
1
2
1
3
2
3
@@
We identify $\mathscr{B}^{k}$ and $\mathscr{T}^{k}$ by the one-to-one correspondence where $i_{\alpha}=0,1$ in ${\bf i}\in\mathscr{B}^{k}$ is regarded as the multiplicity of the letter $\alpha$ in ${\bf I}\in\mathscr{T}^{k}$. The arrays $\bf i$} in (3.1) and ${\bf I}\in\mathscr{T}^{k}$ in (3.4) will be referred to as the multiplicity representation and the tableau representation, respectively.
3.2. $S^{k,1}(z)$ from antisymmetric fusion
Here we present the matrix $S^{k,1}(z)\in\mathrm{End}(V^{k}\otimes V^{1})$ with $0\leq k\leq n+1$ satisfying the Yang-Baxter relation. For simplicity, we shall often suppress the superscripts and denote it by $S(z)$. Its action on the base vectors of $V^{k}\otimes V^{1}$ is given by
$S(z)(v_{\bf i}\otimes v_{{\bf e}_{j}})$ $=\sum_{{\bf a}\in\mathscr{B}^{k},{\bf e}_{b}\in\mathscr{B}^{1}}S(z)^{{\bf a},{\bf e}_{b}}_{{\bf i},{\bf e}_{j}}\,v_{\bf a}\otimes v_{{\bf e}_{b}}\qquad({\bf i}\in\mathscr{B}^{k},{\bf e}_{j}\in\mathscr{B}^{1}).$ (3.5)
The matrix element will be depicted as in Figure 2.
[…]
Figure 2. A typical matrix element.
@@ 3.2. $S^{k,1}(z)$ from antisymmetric fusion
The $R$-matrix $S^{k,1}(z)$ can be constructed in two ways. A traditional method is the antisymmetric fusion of the $k=1$ case $S^{1,1}(z)$. We will explain it in this subsection. Another method, which is less known but actually much simpler and efficient, is based on 3D integrability. It is summarized in Appendix B.
We write the elements $S(z)^{{\bf e}_{a},{\bf e}_{b}}_{{\bf e}_{i},{\bf e}_{i}}$ of $S^{1,1}(z)$ simply as $S(z)^{a,b}_{i,j}$. They are zero except the following:
$S(z)^{i,i}_{i,i}=1-tz,\quad S(z)^{i,j}_{i,j}=(1-z)t^{[i>j]}\;(i\neq j),\quad S(z)^{j,i}_{i,j}=(1-t)z^{[i<j]}\;(i\neq j).$ (3.6)
They will be depicted as Figure 3.
[…]
Figure 3. Nonzero elements of $S^{1,1}(z)$ with $0\leq i\neq j\leq n$. Horizontal arrows, corresponding to $k=1$, are depicted as ordinary thin ones.
Recall that $V^{1}$ has been identified with the space of local states of the $t$-PushTASEP ${\mathsf{V}}=\bigoplus_{i=0}^{n}\mathbb{C}{\mathsf{v}}_{i}$ as in (3.3). Then the above definition is rephrased as
$S(z)=\sum_{a,b,i,j=0}^{n}S(z)^{a,b}_{i,j}E_{ai}\otimes E_{bj}\in\mathrm{End}({\mathsf{V}}\otimes{\mathsf{V}}),$
where $E_{ij}{\mathsf{v}}_{l}=\delta_{j,l}{\mathsf{v}}_{i}$. It satisfies the Yang-Baxter equation
$S_{1,2}(x)S_{1,3}(y)S_{2,3}(y/x)=S_{2,3}(y/x)S_{1,3}(y)S_{1,2}(x).$
Let $P(\mathsf{u}\otimes\mathsf{v})=\mathsf{v}\otimes\mathsf{u}$ be the transposition. From (3.6), one sees that the image
$\mathrm{Im}PS(t^{-1})=\bigoplus_{0\leq i<j\leq n}\mathbb{C}({\mathsf{v}}_{i}\otimes{\mathsf{v}}_{j}-{\mathsf{v}}_{j}\otimes{\mathsf{v}}_{i})$
is the space of antisymmetric tensors. The Yang-Baxter equation multiplied with $P_{1,2}$ from the left reads $P_{1,2}S_{1,2}(x)S_{1,3}(y)S_{2,3}(y/x)=S_{1,3}(y/x)S_{2,3}(y)P_{1,2}S_{1,2}(x)$. The choice $x=t^{-1}$ here implies that
$S_{1,3}(ty)S_{2,3}(y)\in\mathrm{End}({\mathsf{V}}\otimes{\mathsf{V}}\otimes{\mathsf{V}})$
preserves the space
$\left(\mathrm{Im}P_{1,2}S_{1,2}(t^{-1})\right)\otimes{\mathsf{V}}=\left(\bigoplus_{0\leq i<j\leq n}\mathbb{C}({\mathsf{v}}_{i}\otimes{\mathsf{v}}_{j}-{\mathsf{v}}_{j}\otimes{\mathsf{v}}_{i})\right)\otimes{\mathsf{V}}\subset{\mathsf{V}}\otimes{\mathsf{V}}\otimes{\mathsf{V}}.$
Consider the following operator with an overall scalar factor $d_{k}(z)^{-1}$ which is included to validate the forthcoming Theorem 3.1:
$d_{k}(z)^{-1}S_{1,k+1}(z)S_{2,k+1}(t^{-1}z)\cdots S_{k,k+1}(t^{-k+1}z)\in\mathrm{End}({\mathsf{V}}^{\otimes k}\otimes{\mathsf{V}})\quad(0\leq k\leq n+1),$ (3.7)
$d_{k}(z)=(1-t^{-1}z)(1-t^{-2}z)\cdots(1-t^{-k+1}z).$ (3.8)
By extending the above argument, one can show that (3.7) can be restricted to $\bigwedge^{k}\otimes{\mathsf{V}}$, where $\bigwedge^{k}$ is the subspace of ${\mathsf{V}}^{\otimes k}$ spanned by the degree $k$ antisymmetric tensors
$\bigwedge\nolimits^{\!k}$ $=\bigoplus_{{\bf I}\in\mathscr{T}^{k}}\mathbb{C}\,\nu_{\bf I},\qquad\nu_{\bf I}=\sum_{\sigma\in\mathfrak{S}_{k}}\mathrm{sgn}(\sigma){\mathsf{v}}_{I_{\sigma(k)}}\otimes\cdots\otimes{\mathsf{v}}_{I_{\sigma(1)}}\in{\mathsf{V}}^{\otimes k},$ (3.9)
where $\mathfrak{S}_{k}$ is the symmetric group of degree $k$, and $\mathrm{sgn}(\sigma)$ denotes the signature of the permutation $\sigma$. The set $\mathscr{T}^{k}$ has been defined in (3.4). We identify $\bigwedge^{k}$ in (3.9) with $V^{k}$ in (3.2) via
$V^{k}\ni v_{\bf i}=\nu_{\bf I}\in\bigwedge\nolimits^{\!k},\quad\mathscr{B}^{k}\ni{\bf i}={\bf I}\in\mathscr{T}^{k},$ (3.10)
where the bijective correspondence between the multiplicity arrays in $\mathscr{B}^{k}$ (3.1) and the column strict tableaux in $\mathscr{T}^{k}$ has been explained after (3.4).
The $R$-matrix $S^{k,1}(z)$ is obtained as the restriction of the operator (3.7) to $\bigwedge^{k}\otimes{\mathsf{V}}$ according to the above identification. This construction leads to the following formula for its matrix elements:
$S(z)^{{\bf a},{\bf e}_{b}}_{{\bf i},{\bf e}_{j}}=d_{k}(z)^{-1}\sum_{b_{1},\ldots,b_{k-1}\in\{0,\ldots,n\}}\sum_{\sigma\in\mathfrak{S}_{k}}\mathrm{sgn}(\sigma)\\ \times S(z)^{A_{k},b}_{I_{\sigma(k)},b_{k-1}}S(t^{-1}z)^{A_{k-1},b_{k-1}}_{I_{\sigma(k-1)},b_{k-2}}\cdots S(t^{-k+1}z)^{A_{1},b_{1}}_{I_{\sigma(1)},j},$ (3.11)
where $b,j\in\{0,\ldots,n\}$ and ${\bf a},{\bf i}\in\mathscr{B}^{k}$ are multiplicity arrays. The arrays $(A_{1},\ldots,A_{k})$ and $(I_{1},\ldots,I_{k})\in\mathscr{T}^{k}$ (3.4) are the tableau representations of $\bf a$} and ${\bf i}\in\mathscr{B}^{k}$, respectively. We note that the sign factor $\mathrm{sgn}(\sigma)$ is a simplifying feature of the current gauge, in contrast to the $t$-dependent factor $(-t)^{\mathrm{length}(\sigma)}$, which is commonly encountered in the conventional gauge. To summarize, the weight is given as in Figure 4.
[…]
Figure 4. The weight of the element in (3.11), where $b_{r}$ is assigned to the vertical edge between $A_{r}$ and $A_{r+1}$. Note that $I_{1}<I_{2}<\dots<I_{k}$ and $A_{1}<A_{2}<\dots<A_{k}$ by (3.4).
We use the notation $i,j]=\{i,i+1,\dots,j-1,j\}$ for $i\leq j$.
Theorem 3.1.
$S(z)^{{\bf a},{\bf e}_{b}}_{{\bf i},{\bf e}_{j}}$ $=\delta^{{\bf a}+{\bf e}_{b}}_{{\bf i}+{\bf e}_{j}}(-1)^{a_{0}+\cdots+a_{j-1}+i_{0}+\cdots+i_{b-1}}t^{a_{j+1}+\cdots+a_{n}}(1-t^{a_{j}}z^{\delta_{b,j}})z^{[j>b]}.$ (3.12)
We include a proof in Appendix A. The sign factor in (3.12) can also be expressed as $(-1)^{a_{r+1}+\cdots+a_{s-1}}$ with $r=\min(j,b)$ and $s=\max(j,b)$. The property that $S(z)^{{\bf a},{\bf e}_{b}}_{{\bf i},{\bf e}_{j}}=0$ unless ${\bf a}+{\bf e}_{b}={\bf i}+{\bf e}_{j}$ is referred to as weight conservation. Matrix elements (3.12) have appeared in a different gauge as the basic ingredient in the vertex operator approach in [DO94]. The basic case $k=1$ in (3.6) are positive in some range of $t,z$, and the sum $\sum_{0\leq a,b\leq n}S(z)^{a,b}_{i,j}=1-tz$ is independent of $i,j$. It is well known that these properties can be utilized to construct a Markov Matrix of multispecies ASEP; see Section 7.
On the other hand for $k\neq 1$ in general, (3.12) is neither positive definite nor negative definite for fixed $t$ and $z$. Furthermore, the summation $\sum_{{\bf a}\in\mathscr{B}^{k},{\bf e}_{b}\in\mathscr{B}^{1}}S(z)^{{\bf a},{\bf e}_{b}}_{{\bf i},{\bf e}_{j}}$ does not become independent of ${\bf i},{\bf e}_{j}$. Consequently, the $R$-matrix $S^{k,1}(z)$ is not stochastic by itself in the sense of [KMMO16]. What is intriguing, as revealed by our subsequent analysis, is that the Markov Matrix of the multispecies $t$-PushTASEP is nonetheless reproduced as a suitable linear combination of the transfer matrices constructed from $S^{0,1}(z),S^{1,1}(z),\ldots,S^{n+1,1}(z)$.
The matrix $S^{k,1}(z)$ given above is member of a larger family of $R$-matrices $S^{k,l}(z)\in\mathrm{End}(V^{k}\otimes V^{l})$ with $0\leq k,l\leq n+1$, and they satisfy the Yang-Baxter equation
$S^{k_{1},k_{2}}_{1,2}(x)S^{k_{1},k_{3}}_{1,3}(xy)S^{k_{2},k_{3}}_{2,3}(y)=S^{k_{2},k_{3}}_{2,3}(y)S^{k_{1},k_{3}}_{1,3}(xy)S^{k_{1},k_{2}}_{1,2}(x).$ (3.13)
An efficient method to construct $S^{k,1}(z)$ by 3D integrability is explained in Appendix B.